# Investigating Language Preference of Multilingual RAG Systems

### v2

@@ G.1 Failure Case
MLR
We present a failure case of the MLR metric in Table 13. Due to the difficulty of translating documents in low-resource languages, repetitive phrases such as Changing the line-up appear in the translated passage. This repetition causes the re-ranker to misinterpret the content, leading to an improvement in the rank even though the content is irrelevant.
@@
Table 13: A failure case of MLR because bad translation quality due to difficulty in translating low-resource language.
@@ Appendix H Language Preference of Generators in average
Appendix I Ablation study of DKM-RAG
To prove the effectiveness of concatenating translated passages and refined passages in DKM-RAG framework, we provide an ablation study of each component in DKM-RAG. As stated in Table 15, removing any component from DKM-RAG results in decreased performance, highlighting that every part is crucial to its overall effectiveness.
DKM-RAG w/o $P_{\text{refined}}$ w/o $P_{\text{translated}}$
$L_{q}=$ en
aya-expanse-8b 82.60 79.34 81.10
Phi-4 82.59 78.89 81.08
Qwen2.5-7B-Instruct 82.60 79.11 81.06
Llama3.1-8B-Instruct 82.57 79.28 81.19
$L_{q}=$ zh
aya-expanse-8b 44.57 38.31 39.44
Phi-4 44.56 36.76 38.95
Qwen2.5-7B-Instruct 44.70 38.31 39.78
Llama3.1-8B-Instruct 44.51 38.48 39.35
$L_{q}=$ ko
aya-expanse-8b 55.01 49.66 46.15
Phi-4 54.82 49.25 45.24
Qwen2.5-7B-Instruct 54.85 49.44 45.32
Llama3.1-8B-Instruct 54.99 49.87 45.55
Appendix J MLR Analysis
We prove the effectiveness of our proposed language preference metric, MLR by comparing language preference between MLR score and the average document language ratio of retrieved documents for each dataset. As stated in Table 16, the tendency of average language ratio of retrieved documents and MLR score is similar. To prove it, we also report Pearson and Spearman correlation coefficients and each p-value between them. Pearson value (0.98558) indicates a very strong positive linear correlation between the average MKQA language distribution values (mkqa_avg) and the MLR (Preference) scores. The p-value (7.75e-10) is extremely small, showing that the probability of observing such a strong correlation by chance is almost negligible. In short, there is a statistically significant, nearly perfect linear relationship between these two sets of values. Similarly, the Spearman value (0.86264) also indicates a strong association, and the corresponding p-value (1.47e-4) confirms that this correlation is statistically significant. By these results, we prove that MLR is efficient for measuring language preference of retriever.
@@ Appendix I Ablation study of DKM-RAG
Table 16: Language distribution ratios of documents retrieved from datasets composed of each query language. The table lists the raw MKQA language distribution values (without the percent sign) for each dataset. The row mkqa_avg shows the average distribution across all MKQA datasets for each language, while the row MLR (Preference) provides the corresponding MLR scores. Additionally, we report Pearson and Spearman correlation coefficients between MLR and mkqa_avg.

### v3

@@ G.1 Failure Case
MLRS
We present a failure case of the MLRS metric in Table 15. Due to the difficulty of translating documents in low-resource languages, repetitive phrases such as Changing the line-up appear in the translated passage. This repetition causes the re-ranker to misinterpret the content, leading to an improvement in the rank even though the content is irrelevant.
@@
Table 15: A failure case of MLRS because bad translation quality due to difficulty in translating low-resource language.
@@ Appendix H Language Preference of Generators in average
Appendix I MLRS Analysis
We prove the effectiveness of our proposed language preference metric, MLRS by comparing language preference between MLRS score and the average document language ratio of retrieved documents for each dataset. As stated in Table 17, the tendency of average language ratio of retrieved documents and MLRS score is similar. To prove it, we also report Pearson and Spearman correlation coefficients and each p-value between them. Pearson value (0.98558) indicates a very strong positive linear correlation between the average MKQA language distribution values (mkqa_avg) and the MLRS (Preference) scores. The p-value (7.75e-10) is extremely small, showing that the probability of observing such a strong correlation by chance is almost negligible. In short, there is a statistically significant, nearly perfect linear relationship between these two sets of values. Similarly, the Spearman value (0.86264) also indicates a strong association, and the corresponding p-value (1.47e-4) confirms that this correlation is statistically significant. By these results, we prove that MLRS is efficient for measuring language preference of retriever.
@@ Appendix I MLRS Analysis
Table 17: Language distribution ratios of documents retrieved from datasets composed of each query language. The table lists the raw MKQA language distribution values (without the percent sign) for each dataset. The row mkqa_avg shows the average distribution across all MKQA datasets for each language, while the row MLRS (Preference) provides the corresponding MLRS scores. Additionally, we report Pearson and Spearman correlation coefficients between MLRS and mkqa_avg.